OpenStudy (anonymous):

Let f(x) = 4x − 3. If f(a) = 9 and f(b) = 5, then what is f(a + b) ?

OpenStudy (anonymous):

What you are going to want to do with this problem is add up (4(9) -3) + (4(5) - 3) Do you understand how I substituted those values?

OpenStudy (anonymous):

No I don't, if you don't mind could you explain it?

OpenStudy (anonymous):

Because the problem requires f(a) + f(b), I added up the equation with the value of f(a) and f(b) together. f(a) = 4(9) - 3 f(b) = 4(5) - 3 I simply added the two equations together. I got the values from the equations given.

OpenStudy (whpalmer4):

no, it requires \(f(a+b)\ which is not necessarily the same as \(f(a)+f(b)\!

OpenStudy (anonymous):

True whpalmer, so regard that note. Besides the mistake the equation is still set up correctly. I hope that doesn't confuse you too much.

OpenStudy (cwrw238):

no kirby f(a)= 9 means 4*a - 3 = 9 making a = 3

OpenStudy (whpalmer4):

if \(f(x) = 4x-3\) and \(f(a) = 9\) then \[f(a) = 4a-3 = 9\]solving for \(a\) \[4a-3=9\]\[4a=12\]\[a=3\] Now do the same for \(b\), then add \(a+b\) and plug it into \(f(x) = 4x-3\)

OpenStudy (cwrw238):

b = 2 so we need to find f(2+3) = f(5)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Notice that \(x\mapsto4x-3\) is an affine map so we can see that \(f(a+b)=f(a)+f(b)+3\) ergo \(f(a+b)=9+5+3=17\)

OpenStudy (anonymous):

no need to even solve for \(a,b\) just reason it out...$$f(a+b)=4(a+b)-3=(4a-3)+(4b-3)+3=f(a)+f(b)+3$$

OpenStudy (whpalmer4):

true enough, but someone asking for help with this problem is unlikely to know about affine maps, and likely to draw the wrong conclusion that it is safe to assume that f(a+b) = f(a) + f(b) in the future

OpenStudy (anonymous):

nobody said \(f(a+b)=f(a)+f(b)\) but if \(f(x)=kx\) it's rather obvious that \(f(a+b)=f(a)+f(b)\)... so naturally if we shift \(f\) vertically by \(h\) you can intuitively see why \(f(a+b)-h=f(a)-h+f(b)-h\) ergo \(f(a+b)=f(a)+f(b)-h\)

OpenStudy (whpalmer4):

actually, if you re-read some of the first responses, you'll see that someone did.

OpenStudy (anonymous):

alright well I meant in this current conversation... I pay no attention to wrong answers

OpenStudy (anonymous):

:-p

OpenStudy (whpalmer4):

if only we could assume the original poster had the same filter, our life would be much easier :-)

OpenStudy (whpalmer4):

anyhow, thanks for providing me with the "learn something new every day" moment (hadn't heard of affine maps) :-)